Survey of RepRap and Makerbot Community Members

script for internet survey

- Erik, Jeroen and Eric - version 23 February 2010

Preface

- This survey is about your motives and efforts to use and further develop the RepRap, or any derivate machine (e.g., Makerbot). We want to better understand conditions under which open development of physical products works. We also want to learn how it differs from proprietary modes of development.

- It is part of a research program initiated by the MIT Sloan school of management, (Cambridge, USA) and Tilburg University (the Netherlands) to study open innovation by online communities. I will use the data to write my master thesis.

- On average, the survey will take 15 to 20 minutes.

- All information received will remain ANONYMOUS and be treated strictly CONFIDENTIAL.

- Do you have any questions or comments? Please contact me at: phone +31(0)137113076, e-mail info@erikdebruijn.nl.

- Many thanks for your co-operation!

Best wishes,

Erik de Bruijn

SECTION A: TYPE OF USER

Intro1 The first questions are about the kind of machine(s) that you use and/or are building.

A1 Please mark what kind of machine(s) you currently use or are building. (Multiple answers possible)

I am having one in operation I am currently building one

a. Bits From Bytes Rep(st)Rap v 3.0 (acrylic)
b. Bits From Bytes Rep(st)Rap v 2.0 (acrylic)
c. Bits From Bytes Rep(st)Rap (earlier version)
d. RepRap Darwin (built from RP parts)
e. RepRap Mendel (built from RP parts)
f. Makerbot

g. McWire RepStrap
h. Other RepStrap
i. My own RepStrap/3D printer design

j.

Existing CNC-Mill with RepRap extruder toolhead (or using an off-the-shelf cartesian bot)

k. Other machine

A2 When did you start building your first machine? . . . . . . . . . . (month) . . . . . . .. (year)

A3 Do you have or build your own machine, or do you share one with others? (Multiple answers possible) I have my own machine I am sharing a machine with others

A4 (if A3=2) With how many others do you share your machine? . . . . . . .others

If A1a-A1k are only ‘currently building one’ then Go to A6

A5 Please mark how often you engage in the following activities.

never  barely sometimes often always
a. Building the machine; getting it to work

b.

Using a machine to print objects for yourself (e.g., rapid prototyping, direct part production, presentation models, etc)

c.

Developing improvements for the machine (e.g., hardware, software)

d.

Helping other users (e.g., technical assistance, sending parts, documenting your work on the Web, etc)?

e.

Improving your skills (e.g., learning electronics to better work with and/or improve the machine).

A6 How much time do you weekly spend on working with or developing your machine? (please estimate - including everything, e.g. all invested time to build, use and improve the

machine, and or help other users) . . . . . . .hours per week

A7 (if A6>0) How is this time distributed across…

a. Building the machine; getting it to work?  . . . ..
b. Printing objects?  . . . ..
c. Developing improvements for the machine (in order to print what I need)?  . . . ..
d. Developing improvements for the machine (just to make it better)?

e.

Helping other users (e.g., technical assistance, sending parts, documenting your work on the Web, etc)?

. . . ..

f. Improving your skills?  . . . ..

100%

A8 (if A7d > 0) How do you help other users (multiple answers possible)? By meeting people to provide technical assistance or advice By giving instructions or answering questions via the Web, e-mail or phone By sending tool or parts By posting information on the Web By improving or extending information on the Web Other, please specify. . . . . . . . . . . . . . . . . . ..

A9 How much money did you spend on the machine since you started with it? (please estimate – if your local currency is unavailable, please convert to Euros or US dollars) . . . . . . . . . ..Euros/dollars/yen/UK pounds

A11 In the past three months, how many other users of 3D printers did you communicate with?

(personal communication by means of e-mail, telephone, face to face, etc – please estimate) ….users

A12 (if A11>0) And in the past three months, how many other users did you meet in person? ….users

A13 Are you member of a local community of 3D printer users, i.e. that you meet in person from

time to time?

yes no

A14 (if A13=yes) How many members does this local community have? ….members

A15 (if A13=yes) Does this local community have a Web presence (e.g., website, wiki, blog)? yes no

A16 (if A15=yes) Can you give us the URL of your local community’s website?

. . . . . . . . . . . . . . . . . . . . . . . . . . .

A17 What motivates you to use or build your machine? Please distribute 100 points according to

what you find most important.

a. Building the machine; getting it to work  . . . ..
b. Printing objects  . . . ..
c. Developing improvements for the machine (in order to print what I need)  . . . ..
d. Developing improvements for the machine (just to make it better)

e.

Helping other users (e.g., technical assistance, sending parts, documenting your work on the Web, etc)

. . . ..

f. Improving your skills  . . . ..

100%

A18 (if A17f > 0 ) What other reasons do you have to use or build the machine? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

SECTION B: ADOPTING THE MACHINE

Intro2 The following questions are about the first machine that you built (or are still building).

B1 How long did it take you (so far) to get this machine working? ….hours/days/weeks/months

B2 Did you face any problems to get this machine working? yes no

B3 (if B2 = yes) Where these problems primarily related to the hardware, the software, or both? hardware software both hardware and software

B4 (if B3=1 or 3) What sources did you consult to learn about any hardware-related problems? (multiple answers possible) Online wiki’s, forums and blogs (e.g., RepRap site, BfB, Makerbot) Online video sharing sites (e.g., Youtube) Personal communication with other users (e-mailing, chatting, telephone calls) Personal assistance by other users (clubs, one-to-one visits) Physical inspections of other machines Other, please specify. . . . . . . . . . . . . . . . . .

B5 (if B3 = 1 or 3) Did you manage to solve your hardware-related problems? yes, completely yes, partially no or not yet

B6 (if B3=2 or 3) What sources did you consult to learn about any software-related problems? (multiple answers possible) Online wiki’s, forums and blogs (e.g., RepRap site, BfB, Makerbot) Online video sharing sites (e.g., Youtube) Personal communication with other users (e-mailing, chatting, telephone calls) Personal assistance by other users (clubs, one-to-one visits) Physical inspections of other machines Other, please specify. . . . . . . . . . . . . . . . . .

B7 (if B3 = 2 or 3) Did you manage to solve your software-related problems? yes, completely yes, partially no or not yet

SECTION C: INNOVATING THE SOFTWARE

Intro3 The following questions relate to any software that you may develop or improve for the machine.

C1 Did you ever modify or develop any NEW software for the machine? (by programming original

code) Yes No

If C1=no Go to Intro4

C2 Please describe your most important software modification/creation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

C3 The next questions are concerned with this specific piece of software. Please mark if the following statements are true. This software… (multiple answers possible)

… added new functions to the machine.
… improved the convenience of the machine.

… increased the performance of the machine (e.g., speed, resolution).

C3a Were you the first to modify or develop this software? I think that… (Please mark) … I was first … others were developing/modifying similar things

C4 What motivated you to develop this software? Please distribute 100 points according to what

motivated you most.

a. I needed this software to build the machine, i.e. to get it to work  . . . ..
b. I needed it to print specific objects  . . . ..
c. I enjoy developing improvements for the machine  . . . ..
d. I wanted to help other users  . . . ..
e. I wanted to improve my skills  . . . ..
f. Any other reason  . . . ..

100%

C5 (if C4f > 0 ) For what other reasons did you develop this software? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

C6 What or who did you consult to develop this software? (multiple answers possible) Online wiki’s, forums and blogs (e.g., RepRap site, BfB, Makerbot) Online video sharing sites (e.g., Youtube) Personal communication with other users (e-mailing, chatting, telephone calls) Personal assistance by other users (clubs, one-to-one visits) Physical inspections of other machines Other, please specify. . . . . . . . . . . . . . . . . .

C7 Did you collaborate with other users to develop this software? yes no

C8 (if C7=yes) Can you estimate how many others were involved? . . . ..other users

C9 (if C7=yes AND A13=yes) Were these collaborators members of your local community? yes partially no C10 Can you estimate how much time you spent on developing/modifying this software? (Please estimate) . . . . . . . . . ..hours/days/weeks/months

C11 Did you spend any money to develop this software? yes no

C12 (if C11=yes) Can you estimate how much? (Please estimate – if your local currency is unavailable, please convert to Euros or US dollars) . . . . . . . . . .UK Pounds/dollars/yen/euros/ etc

C13 Did you reveal the details of this software to other people? yes no

C14 (if C13=yes) How did you reveal it? (multiple answers possible) Freely revealed it on the Web (e.g., Wiki, Forum, Blog) Selectively revealed it to specific other users (e.g., e-mail, messaging, etc) Otherwise, please specify. . . . . . . . . . . . . . . . . .

C15 (if C13=yes) To the best of your knowledge, have any other persons adopted this software? Yes No

C16 (if C15=yes AND A13=yes) Are these persons all from your local community? Yes No, also people outside my local community adopted it

C17 In general, what would keep other people from adopting this software? (multiple answers possible) It is specific to my needs, not very useful to others It is still very experimental, i.e. needs further development Others find it hard to see the benefit It takes a lot of time to make it work It is hard to properly communicate, i.e. to explain how it should be done It is expensive, i.e. requires a great deal of investment It is difficult to integrate with other parts of the machine Other, please specify. . . . . . . . . . . . . . . . . . None of these, it is easy to adopt this software

SECTION D: INNOVATING THE HARDWARE Intro4 The following questions relate to any hardware that you may develop or improve for the machine.

D1 Did you ever modify or develop any NEW hardware for the machine? Yes No

Skip If D1=no Go to E1

D2 Please describe your most important hardware modification or creation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

D3 The next questions are concerned with this specific piece of hardware. Please mark if the following statements are true. This hardware… (multiple answers possible)

… added new functions to the machine.
… improved the convenience of the machine.
… increased the performance of the machine (e.g., speed, resolution).

D3a Were you the first to modify or develop this hardware? I think that… (Please mark) … I was first … others were developing/modifying similar things

D4 What motivated you to develop this hardware? Please distribute 100 points according to what

motivated you most.

a. I needed this hardware to build the machine, i.e. to get it to work  . . . ..
b. I needed it to print specific objects  . . . ..
c. I enjoy developing improvements for the machine  . . . ..
d. I wanted to help other users  . . . ..
e. I wanted to improve my skills  . . . ..
f. Any other reason  . . . ..

100%

D5 (if D4f > 0 ) For what other reasons did you develop this hardware? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

D6 What or who did you consult to develop this hardware? (multiple answers possible) Online wiki’s, forums and blogs (e.g., RepRap site, BfB, Makerbot) Online video sharing sites (e.g., Youtube) Personal communication with other users (e-mailing, chatting, telephone calls) Personal assistance by other users (clubs, one-to-one visits) Physical inspections of other machines Other, please specify. . . . . . . . . . . . . . . . . .

D7 Did you collaborate with other users to develop this hardware? yes no

D8 (if D7=yes) Can you estimate how many others were involved? . . . ..other users

D9 (if D7=yes AND A13=yes) Were these collaborators members of your local community? yes partially no

D10 Can you estimate how much time you spent on modifying/developing this hardware? (Please estimate) . . . . . . . . . ..hours/days/weeks/months

D11 Did you spend any money to develop or modify this hardware? yes no

D12 (if D11=yes) Can you estimate how much? (Please estimate – if your local currency is unavailable, please convert to Euros or US dollars) . . . . . . . . . .UK Pounds/dollars/yen/euros/ etc

D13 Did you reveal the details of this hardware to other people? yes no

D14 (if D13=yes) How did you reveal it? (multiple answers possible) Freely revealed it on the Web (e.g., Wiki, Forum, Blog) Selectively revealed it to specific other users (e.g., e-mail, messaging, etc) Otherwise, please specify. . . . . . . . . . . . . . . . . .

D15 (if D13=yes)

To the best of your knowledge, have any other persons adopted this software? Yes No

D16 (if D15=yes AND A13=yes) Are these persons all from your local community? Yes No, also people outside my local community adopted it

D17 In general, what would keep other people from adopting this hardware? (multiple answers possible) It is specific to my needs, not very useful to others It is still very experimental, i.e. needs further development Others find it hard to see the benefit It takes a lot of time to make it work It is hard to properly communicate, i.e. to explain how it should be done It is expensive, i.e. requires a great deal of investment It is difficult to integrate with other parts of the machine None of these, it is quit easy to adopt this hardware Other, please specify. . . . . . . . . . . . . . . . . . None of these, it is easy to adopt this hardware

SECTION E: THINGIVERSE

E1 Did you ever print something that you downloaded from Thingiverse?

Yes No E2 (if E1=yes) How many objects have you printed that you got from Thingiverse? (please estimate)

. . . . . . . . . .objects

E3 (if E2>0) For how many of these printed objects have you uploaded a picture on Thingiverse? (please estimate)

. . . . . . . . . objects

E4

Have you ever digitally designed your own physical objects? Yes

No E5 (if E4=yes) How many designs did you make? (please estimate)

. . . . . . . . . designs

E6 (if E4=yes) How many of these designs did you post on Thingiverse? (please estimate)

. . . . . . . . . designs

SECTION F: DEMOGRAPHY AND GENERAL QUESTIONS

Intro5 Finally, we would like you to answer some general questions.

F3 What is your age? ….years

F4 Are you currently employed, self-employed, a student, retired or not working? employed self-employed student retired/not working other, please specify. . . . . . ..

F5 (if F4 = employed or self-employed) Please describe you job or kind of business that you are in. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

F6 (if F4 = student) What kind of study do you do (e.g., engineering, information technology, business administration,

etc) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

F7 What is your highest educational attainment? high school bachelor degree master degree doctorate degree other, please specify. . . . . . . . . . . . ..

F10 How do you assess your skills in the following field of interest? Please mark your level of expertise.

My level of expertise in … is…  rookie  novice  skilled  expert
a. programming original software code
b. tinkering with electronics
c. mechanical systems/machines
d. CAD software
e. CAM, CNC, RP systems and/or tooling
f. design

F11 Are you currently listed on the RepRap Worldmap? (NOT the Makerbot Map) Yes No

F12 Would you like to be notified of our research findings? (We will send our research report to your e-mail address) Yes, send the report to . . . . . . . . . (e-mail address) No, thanks

F13 We would like to invite more users of the RepRap (or derivative machines) to take this survey.

Can you give us the details (name and e-mail adresses) of three other users?

(we will carefully check that no-one is invited twice – all details will be treated confidentially and destroyed after our research)

Yes No, I would rather not give such details

F14 (als F13=yes) Please give us the details of up to three more users. name e-mail . . . .. . . . .. . . . .. . . . .. . . . .. . . . ..

End This is the end of the survey. Many thanks for your co-operation.